OpenStudy (anonymous):

Solve the system by substitution. -x-y-z=-8 -4x+4y+5z=7 2x+2z=4 Please show step by step instructions.

OpenStudy (anonymous):

@Tessa___98 Are these 3 different equations or all together?

OpenStudy (anonymous):

3 different equations

OpenStudy (anonymous):

@Tessa___98 ok lets start with number one

OpenStudy (anonymous):

ook

OpenStudy (anonymous):

so what's first??

OpenStudy (anonymous):

@Tessa___98 1. solve for x -x -y -z = -8 add y to both sides -x -y -z + y = -8 + y -x -z = y - 8 add z to both sides -x -z + z = y - 8 + z -x = y + z - 8 divide both sides by -1 -x/-1 = y + z -8/1 answer -> x = -y - z + 8

OpenStudy (anonymous):

Then what @cutelilcigarette??

OpenStudy (anonymous):

@Tessa___98 2. solving for x - 4x + 4y + 5z = 7 add -4y to both sides -4x + 4y + 5z + -4y = 7 + -4y -4x + 5z = -4y + 7 add -5z to both sides -4x + 5z + -5z = -4y + 7 + -5z -4x = -4y - 5z + 7 divide both sides by -4 -4x/4 = -4y - 5z + 7/-4 answer -> x = y + 5/4z + -7/4

OpenStudy (anonymous):

Thanks a bunch for helping me!!

OpenStudy (anonymous):

@Tessa___98 3. solving for x again 2x + 2z = 4 add -2z to both sides 2x + 2z + -2z = 4 + -2z 2x = -2z + 4 divide both sides by 2 2x/2 = -2z + 4/2 x = - z + 2 <- answer

OpenStudy (anonymous):

and your welcome!

OpenStudy (anonymous):

Ook so is that all that needs to be done?? @cutelilcigarette

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

Ook thanks so so much again for helping!!

OpenStudy (anonymous):

welcome <3